OpenStudy (wintersuntime):

may someone please help me on how to find the inverse for the problem I am about to post on the comments. Thanks

OpenStudy (wintersuntime):

\[f(x)=\sqrt[5]{x-3}\]

OpenStudy (zenmo):

substitute y for f(x)

OpenStudy (wintersuntime):

okay

OpenStudy (zenmo):

what do you have now?

OpenStudy (wintersuntime):

\[y=\sqrt[5]{x-3}\]

OpenStudy (zenmo):

Now, switch x and y (hint: they trade places)

OpenStudy (wintersuntime):

\[x=\sqrt[5]{y-3}\]

OpenStudy (zenmo):

Now, we want to strip the right side to the point of "y" being alone.

OpenStudy (wintersuntime):

so do we add 3?

OpenStudy (zenmo):

to begin, we need to get rid of the square root of 5. Do you know how to get rid of it on the right side, and bring it to the left side?

OpenStudy (wintersuntime):

nope

OpenStudy (zenmo):

we want to start with the outer layers first, since we can't add 3 yet.

OpenStudy (wintersuntime):

ok

OpenStudy (zenmo):

to get rid of square root of 5, u want to multiply it on both sides. Multiplying a square root with a square root cancels out.

OpenStudy (wintersuntime):

okay

OpenStudy (wintersuntime):

what do we multiply it by though?

OpenStudy (wintersuntime):

5?

OpenStudy (zenmo):

\[\sqrt[5]{x}=y-3\]

OpenStudy (zenmo):

Do you understand how to get that part?

OpenStudy (wintersuntime):

not really. How do we know what number to multiply it by?

OpenStudy (zenmo):

@myininaya do you know whats "it" called when u multiply the square root to eliminate \[\sqrt[5]{y-3}\] that makes \[\sqrt[5]{}\]. I forgot what it is called.

OpenStudy (zenmo):

\[\sqrt[5]{x}\]

OpenStudy (wintersuntime):

isn't it to square both sides?

OpenStudy (zenmo):

I think that is what it is called.

OpenStudy (wintersuntime):

ok

OpenStudy (zenmo):

Then, since were one step away from solving the inverse problem. Now you add 3 to the left side, and that is your answer.

OpenStudy (zenmo):

Your answer should be \[y=\sqrt[5]{x}+3\]

OpenStudy (wintersuntime):

oh okay thank you so much

OpenStudy (zenmo):

You understand why we don't include "+3" in the square root?

OpenStudy (zenmo):

A slight mistake, people could make.

OpenStudy (wintersuntime):

wait why can't we include add 3 in the square root?

myininaya (myininaya):

I think you meant to raise both sides to the 5th power earlier

myininaya (myininaya):

\[x=\sqrt[5]{y-3} \\ x^5=y-3\]

OpenStudy (zenmo):

Looking at the original equation, 3 is "packaged" together with the "x." Since, we have stripped the outer layer, the square root, \[\sqrt[5]{y-3}\] simply becomes y-3 (since we squared both sides, x becoming \[\sqrt[5]{x}\]

myininaya (myininaya):

then last step add 3 on both sides

OpenStudy (wintersuntime):

so then the answer would be \[x^5+3 \]

OpenStudy (zenmo):

Oops, myininaya is correct. Sorry, I should of done this problem before typing it out. Haha :)

OpenStudy (wintersuntime):

That's okay (:

OpenStudy (zenmo):

but the same case still applies, once you strip the outer layer (square root etc) the 3 becomes free from the y and doesn't become a package with the "y."

OpenStudy (wintersuntime):

okay thanks guys.